OpenStudy (anonymous):

x - 1 4 x # x + 3 x - 1

OpenStudy (anonymous):

i can't understand your problem given that?

OpenStudy (hba):

\[4 x \neq x + 3 \]

OpenStudy (hba):

@NEVALBRUNSON Beta Respond ?

OpenStudy (anonymous):

i trying to post again

OpenStudy (hba):

Okay :)

OpenStudy (anonymous):

x-1/4x*x+3/x-1

OpenStudy (callisto):

You can use the edit question button to edit the question :\

OpenStudy (dls):

\[\LARGE \frac{x-1}{4x} \times \frac {x+3}{x-1} \] is it?

OpenStudy (anonymous):

correct

OpenStudy (dls):

The term (x-1) will cancel out..u will get: \[\LARGE \frac{x+3}{4x}\]

OpenStudy (anonymous):

that is the right answer I am just looking for a place I could go to to do my homework assignments

OpenStudy (anonymous):

dls can we connect